OpenStudy (thedj4jc):

Choose the product. a3b2 · 4ab3 4a3b5 4a3b6 4a4b5

OpenStudy (freckles):

You know exponents are different then coefficients right?

OpenStudy (freckles):

this is what I think you mean: \[a^3b^2 \cdot 4ab^3\] but please correct me if you really do mean \[a3b2 \cdot 4ab3\]

OpenStudy (thedj4jc):

It is!! I keep forgetting it doesn't put them up when I paste

OpenStudy (freckles):

let's look at just the a's for a sec \[a^3 \cdot a \\ a^3 \cdot a^1\] the law here says if you are multiplying with same base add exponents

OpenStudy (freckles):

so what is the product of a^3 and a^1 ?

OpenStudy (thedj4jc):

a^4?

OpenStudy (freckles):

yes you really don't need to do any further work

OpenStudy (freckles):

you listed 3 choices

OpenStudy (freckles):

and only one of those choices has that factor

OpenStudy (thedj4jc):

I see it now!! Thank you so much!!!!

OpenStudy (freckles):

np

OpenStudy (freckles):

but just for fun do you know the product of b^2 and b^3?

OpenStudy (thedj4jc):

Yep!! b^5 :-)

OpenStudy (freckles):

which is also what they have in that third choice :)